# references.bib

### v1

@@ Abstract
The chromatic symmetric $X_{G}$ function is a symmetric function generalization of the chromatic polynomial of a graph, introduced by Stanley (1995). $X_{G}$ is defined by assigning a variable to each color and taking the sum over all proper vertex colorings of $G$ of the product of all variables corresponding to colors used, counted with multiplicity. Stanley (1995) gave an expansion formula for $X_{G}$ in terms of the power sum symmetric functions $p_{\lambda}$ using the principle of inclusion-exclusion. Crew, Pechenik, and Spirkl (2023) defined the Kromatic symmetric function $\overline{X}_{G}$ as a $K$-theoretic analogue of $X_{G}$, constructed in the same way except that each vertex is assigned a nonempty set of colors such that adjacent vertices have nonoverlapping color sets. They defined a $K$-analogue $\overline{p}_{\lambda}$ of the power sum basis and computed the first few coefficients of the $\overline{p}$-expansion of $\overline{X}_{G}$ for some small graphs $G$. They conjectured that the $\overline{p}$-expansion always has integer coefficients and asked whether there is an explicit formula for these coefficients. We give a formula for the $\overline{p}$-expansion of $\overline{X}_{G}$, and prove that the coefficients are indeed always integers.
@@ 1 Introduction
The chromatic symmetric function $X_{G}$ was introduced by Stanley in [stanley1995symmetric] as a symmetric function generalization of the chromatic polynomial [birkhoff1913reducibility]. Stanley gave expansion formulas for $X_{G}$ in a few different bases, including the power sum symmetric functions $p_{\lambda}.$ His classic $p$-expansion formula is an alternating sum over all subsets of the edges of $G$, but he also gave an alternative formulation in terms of the broken circuit complex of $G$, which unlike his subset sum expansion does not involve extra cancellation. Bernardi and Nadeau ([bernardi2020combinatorial]) gave an alternative interpretation of the $p$-expansion of $X_{G}$ in terms of the source components of its acyclic orientations.
@@
One natural question about $\overline{X}_{G}$, posed by the authors of [kromatic2023], is whether there is a $K$-analogue of Stanley’s power sum expansion formula in terms of an appropriate $K$-analogue of the $p$-basis. The authors of [kromatic2023] defined such a $K$-analogue of the $p$-basis, which they called the $\overline{p}$-basis. They gave a table listing the first few coefficients of the $\overline{p}$-expansions of $\overline{X}_{G}$ for several small graphs $G$, which they computed using Sage [sage]. The leading terms (i.e. the terms where $|\lambda|$ is minimal), must always match the corresponding terms from ordinary $p$-expansion of the ordinary chromatic symmetric function $X_{G}.$ (The leading terms are those where $|\lambda|=|V(G)|$ for an unweighted graph, or where $|\lambda|=\omega(G):=\sum_{v\in V(G)}\omega(v)$ for a weighted graph with weight function $\omega:V(G)\to\mathbb{N}$.)
@@
The authors of [kromatic2023] asked whether there is an explicit formula for the $\overline{p}$-expansion of $\overline{X}_{G}$. Here, we give such a formula that applies to any weighted graph $G$ with a vertex weight function $\omega:V\to\mathbb{N},$ in terms of its independence polynomial
$I_{(G,\omega)}(t):=\sum_{S\subseteq V(G)\text{ an independent set}}t^{\omega(S)},$
@@
$\overline{X}_{(G,\omega)}=\sum_{W\subseteq V(G)}(-1)^{|V(G)|-|W|}\prod_{k\geq 1}(1+\overline{p}_{k})^{a_{W}(k)},$
where the exponents $a_{W}(k)$ are defined to satisfy
@@
This idea for where the $(1+t^{k})$ terms come from is the factorization $1+\overline{p}_{k}=\prod_{i\geq 1}(1+x_{i}^{k}).$ As a consequence of Theorem 1, we can write down an explicit formula for the coefficients $[\overline{p}_{\lambda}]\overline{X}_{(G,\omega)}$:
Corollary 2.
@@
$[\overline{p}_{\lambda}]\overline{X}_{(G,\omega)}=\sum_{W\subseteq V(G)}(-1)^{|V(G)|-|W|}\prod_{k=1}^{\ell}\binom{a_{W}(k)}{i_{k}},$
@@
The numbers $a_{W}(k)$ above can be computed recursively as follows:
Proposition 3.
@@
$a_{W}(k)=[t^{k}]\log(I_{(G|_{W},\omega)}(t))+\sum_{d\mid k}\frac{(-1)^{k/d}d\cdot a_{W}(d)}{k}.$
We also show that all coefficients in the above expansion are integers:
Theorem 4.
For every partition $\lambda,$ the coefficient $[\overline{p}_{\lambda}]\overline{X}_{G}$ is an integer.
Another question one can ask is a question of $\overline{p}$-positivity. For $G$ an unweighted graph, it can be shown that the sign of $[p_{\lambda}]X_{G}$ is $(-1)^{|\lambda|-\ell(\lambda)}$. From checking some examples, the coefficients $[\overline{p}_{\lambda}]\overline{X}_{G}$ seem to follow the same pattern of signs, which suggests the following conjecture:
Conjecture 5.
For $G$ unweighted, the sign of $[\overline{p}_{\lambda}]\overline{X}_{G}$ is $(-1)^{|\lambda|-\ell(\lambda)}.$
Note that Conjecture 5 does not apply to weighted graphs. For instance, in the $\overline{p}$-expansion of $K_{21}$ above, the coefficient for $\lambda=3$ is $-1$ while $|\lambda|-\ell(\lambda)=3-1=2$ is even.
The remainder of this paper is organized as follows: In §2, we define some necessary terminology, in §3, we revisit the examples above to illustrate Theorem 1, and in §4, we prove Theorem 1, Corollary 2, Proposition 3, and Theorem 4.
@@ 2 Background
A graph $G$ is a set $V(G)$ of vertices together with a set $E(G)$ of edges, where each edge is an unordered pair of vertices. If $uv$ is an edge, we say that the vertices $u$ and $v$ are adjacent. A weighted graph $(G,\omega)$ is a graph $G$ together with a weight function $\omega:V(G)\to\mathbb{N}$ that assigns a positive integer to each vertex. An independent set or stable set is a subset of $V(G)$ in which no two vertices are adjacent. For $W\subseteq V,$ the induced subgraph $G|_{W}$ with vertex set $v$ is the graph with vertex set $V(G|_{W}):=W$ and edge set $E(G|_{W}):=\{uv\in E(G):u,v\in W\}.$ That is, the edges of $G|_{W}$ are precisely the edges of $G$ with both endpoints in $W.$
A symmetric function $f(x_{1},x_{2},\dots)$ is a function that stays the same under any permutation of the variables, i.e. $f(x_{1},x_{2},\dots)=f(x_{\sigma(1)},x_{\sigma(2)},\dots)$ for any permutation $\sigma$ of $\mathbb{N}.$ A proper coloring of $V(G)$ is a function $\alpha:V(G)\to\mathbb{N}$ that assigns a positive integer valued color to each vertex such that $\alpha(u)\neq\alpha(v)$ whenever $uv\in E(G).$ The chromatic symmetric function is
$X_{G}:=\sum_{\alpha}\prod_{v\in V(G)}x_{\alpha(v)}$
@@
A partition $\lambda=\lambda_{1}\lambda_{2}\dots\lambda_{k}$ is a nonincreasing sequence of positive integers $\lambda_{1}\geq\lambda_{2}\geq\dots\geq\lambda_{k},$ and $\lambda_{1},\dots,\lambda_{k}$ are its parts. We can alternatively write $\lambda=\ell^{i_{\ell}}(\ell-1)^{i_{\ell-1}}\dots 2^{i_{2}}1^{i_{1}}$ to denote the partition with $i_{j}$ parts of size $j$ for each $j=1,2,\dots,\ell-1,\ell.$ For a partition $\lambda=\lambda_{1}\lambda_{2}\dots\lambda_{k},$ the power sum symmetric function $p_{\lambda}$ is
$p_{\lambda}:=\prod_{i=1}^{k}\left(\sum_{j\in\mathbb{N}}x_{j}^{\lambda_{i}}\right)=(x_{1}^{\lambda_{1}}+x_{2}^{\lambda_{1}}+\dots)(x_{1}^{\lambda_{2}}+x_{2}^{\lambda_{2}}+\dots)\dots(x_{1}^{\lambda_{k}}+x_{2}^{\lambda_{k}}+\dots).$
Stanley (1995, [stanley1995symmetric]) proved the following $p$-expansion formula for $X_{G}$:
Theorem 6 (Stanley [stanley1995symmetric], Theorem 1.1).
$X_{G}$ satisfies
$X_{G}=\sum_{S\subseteq E(G)}(-1)^{|S|}p_{\lambda(S)},$
where $\lambda(S)$ is the partition whose parts are the sizes of the connected components of the subgraph of $G$ with vertex set $V(G)$ and edge set $S.$
A proper set coloring is a function $\alpha:V(G)\to 2^{\mathbb{N}}\backslash\{\varnothing\}$ that assigns a nonempty set of positive integer valued colors to each vertex such that adjacent vertices have nonoverlapping color sets, i.e. $\alpha(u)\cap\alpha(v)=\varnothing$ whenever $uv\in E(G).$ For a weighted graph $(G,\omega)$, the Kromatic symmetric function is
$\overline{X}_{(G,\omega)}:=\sum_{\alpha}\prod_{v\in V(G)}\prod_{i\in\alpha(v)}x_{i}^{\omega(v)}.$
@@
We will use a few more definitions to prove Theorem 4. The monomial symmetric function $m_{\lambda}$ is
where the sum is taken over all ordered tuples of $k$ distinct indices $n_{1},\dots,n_{k}\in\mathbb{N}$. The augmented monomial symmetric function $\widetilde{m}_{\lambda}$ is
$\widetilde{m}_{\lambda}:=i_{1}!\dots i_{\ell}!\cdot m_{\lambda},$
$\overline{\widetilde{m}}_{\lambda}:=\overline{X}_{K_{\lambda}},$
Theorem 7 (Crew, Pechenik, and Spirkl [kromatic2023], Proposition 3.4).
The $\overline{\widetilde{m}}$-expansion for $\overline{X}_{G}$ is
3 Examples of Theorem 1
@@
The other subsets of $V(G)$ are $\{v\},$$\{w\},$ and $\varnothing$.∅ . In these cases, we get:
@@
Putting this together, Theorem 1 gives
@@
For the seemingly simpler case of a single unweighted edge [Uncaptioned image], there does not seem to be such a simple formula for the coefficients:
Example 9.
@@
The other stable sets are $\{v\},$$\{w\}$, and $\varnothing$,∅ , which have independence polynomials
@@
Putting this together, Theorem 1 gives
@@
again matching the expansion from §1. The sequence of exponents $2,-1,2,-4,6,\dots$ is sequence A038067 in the Online Encyclopedia of Integer Sequences (OEIS) [oeis].

### v2

@@ Abstract
The chromatic symmetric $X_{G}$ function is a symmetric function generalization of the chromatic polynomial of a graph, introduced by Stanley [stanley1995symmetric]. Stanley [stanley1995symmetric] gave an expansion formula for $X_{G}$ in terms of the power sum symmetric functions $p_{\lambda}$ using the principle of inclusion-exclusion, and Bernardi and Nadeau [bernardi2020combinatorial] gave an alternate $p$-expansion for $X_{G}$ in terms of acyclic orientations. Crew, Pechenik, and Spirkl [kromatic2023] defined the Kromatic symmetric function $\overline{X}_{G}$ as a $K$-theoretic analogue of $X_{G}$, constructed in the same way except that each vertex is assigned a nonempty set of colors such that adjacent vertices have nonoverlapping color sets. They defined a $K$-analogue $\overline{p}_{\lambda}$ of the power sum basis and computed the first few coefficients of the $\overline{p}$-expansion of $\overline{X}_{G}$ for some small graphs $G$. They conjectured that the $\overline{p}$-expansion always has integer coefficients and asked whether there is an explicit formula for these coefficients. In this note, we give a formula for the $\overline{p}$-expansion of $\overline{X}_{G}$, show two ways to compute the coefficients recursively (along with examples), and prove that the coefficients are indeed always integers. In a more recent paper [new_p_expansion], we use our formula from this note to give a combinatorial description of the $\overline{p}$-coefficients $[\overline{p}_{\lambda}]\overline{X}_{G}$ and a simple characterization of their signs in the case of unweighted graphs.
@@ 1 Introduction
Let $G$ be a graph with vertex set $V$. The chromatic symmetric function $X_{G}$ was introduced by Stanley in [stanley1995symmetric] as a symmetric function generalization of the chromatic polynomial [birkhoff1913reducibility]. Stanley gave expansion formulas for $X_{G}$ in a few different bases, including the power sum symmetric functions $p_{\lambda}.$ His classic $p$-expansion formula is an alternating sum over all subsets of the edges of $G$, but he also gave an alternative formulation in terms of the broken circuit complex of $G$, which unlike his subset sum expansion does not involve extra cancellation. Bernardi and Nadeau [bernardi2020combinatorial] gave an alternative interpretation of the $p$-expansion of $X_{G}$ in terms of the source components of its acyclic orientations.
@@
One natural question about $\overline{X}_{G}$, posed by the authors of [kromatic2023], is whether there is a $K$-analogue of Stanley’s power sum expansion formula in terms of an appropriate $K$-analogue of the $p$-basis. The authors of [kromatic2023] defined such a $K$-analogue of the $p$-basis, which they called the $\overline{p}$-basis. They gave a table listing the first few coefficients of the $\overline{p}$-expansions of $\overline{X}_{G}$ for several small graphs $G$, which they computed using Sage [sage]. The leading terms (i.e. the terms where $|\lambda|$ is minimal), must always match the corresponding terms from ordinary $p$-expansion of the ordinary chromatic symmetric function $X_{G}.$ (By leading terms, we mean the terms where $|\lambda|=|V|$ for an unweighted graph, or where $|\lambda|=\omega(G):=\sum_{v\in V}\omega(v)$ for a weighted graph with weight function $\omega:V\to\mathbb{Z}_{>0}$, corresponding to each vertex getting one color.)
@@
The authors of [kromatic2023] asked whether there is an explicit formula for the $\overline{p}$-expansion of $\overline{X}_{G}$. Here, we give such a formula that applies to any weighted graph $G$ with a vertex weight function $\omega:V\to\mathbb{Z}_{>0},$ in terms of its independence polynomial
$I_{(G,\omega)}(t):=\sum_{S\subseteq V\text{ an independent set}}t^{\omega(S)},$
@@
$\overline{X}_{(G,\omega)}=\sum_{W\subseteq V}(-1)^{|V\backslash W|}\prod_{k\geq 1}(1+\overline{p}_{k})^{a_{W}(k)},$
where the exponents $a_{W}(k)$ are the unique (sometimes negative) real numbers satisfying
@@
The idea for where the $(1+t^{k})$ terms come from is the factorization $1+\overline{p}_{k}=\prod_{i\geq 1}(1+x_{i}^{k}),$ which essentially lets us reduce to a single variable problem.
As a consequence of Theorem 1.1, we can write down an explicit formula for the coefficients $[\overline{p}_{\lambda}]\overline{X}_{(G,\omega)}$:
Corollary 1.2.
@@
$[\overline{p}_{\lambda}]\overline{X}_{(G,\omega)}=\sum_{W\subseteq V}(-1)^{|V\backslash W|}\prod_{k=1}^{\ell}\binom{a_{W}(k)}{i_{k}},$
@@
The numbers $a_{W}(k)$ above can be computed recursively in two different ways, and the proofs of these recurrences will make clear why they are uniquely determined. The first (more efficient) way is as follows:
Proposition 1.3.
@@
$a_{W}(k)=[t^{k}]\log(I_{(G|_{W},\omega)}(t))+\sum_{d\mid k,\ d<k}\frac{(-1)^{k/d}d\cdot a_{W}(d)}{k}.$
Note that because of the division by $k$ and the non-integrality of the coefficients of $\log(I_{(G|_{W},\omega)}(t)$, it is not immediately clear from Proposition 1.3 whether the $a_{W}(k)$’s are integers. However, our second recursive formula does make it clear that they are integers:
Proposition 1.4.
The coefficients $a_{W}(k)$ also satisfy the alternative recurrence
$a_{W}(k)=[t^{k}]I_{(G|_{W},\omega)}(t)-\sum_{\lambda\vdash k,\ \lambda\neq k}\ \prod_{j=1}^{k-1}\binom{a_{W}(k)}{i_{j}}.$
Hence, $a_{W}(k)$ is an integer for all $W$ and $k.$
An immediate consequence of Proposition 1.4 and Corollary 1.2 is that the $\overline{p}$-coefficients are integers, resolving the question from [kromatic2023]:
Corollary 1.5.
For every partition $\lambda,$ the coefficient $[\overline{p}_{\lambda}]\overline{X}_{(G,\omega)}$ is an integer.
Corollary 1.5 can also be proven in a different way by considering the transition formula between the $\overline{p}$-basis and another basis defined in [kromatic2023], the $\overline{\widetilde{m}}$-basis. Hence, we will also give a short alternative proof of Corollary 1.5 in §4 based on that transition formula.
Another question one can ask is about the signs of the coefficients. For $G$ an unweighted graph, it can be shown that the sign of $[p_{\lambda}]X_{G}$ is $(-1)^{|\lambda|-\ell(\lambda)}$. In a more recent paper [new_p_expansion], we use Theorem 1.1 to give a combinatorial interpretation of the coefficients $[\overline{p}_{\lambda}]\overline{X}_{G},$ and we show that their signs follow the same pattern as the signs of $[p_{\lambda}]X_{G}$, and that the signs of the exponents $a_{W}(k)$ alternate:
Theorem 1.6.
For $G$ unweighted, the sign of $[\overline{p}_{\lambda}]\overline{X}_{G}$ is $(-1)^{|\lambda|-\ell(\lambda)}.$ Also, the sign of $a_{W}(k)$ is $(-1)^{k+1}.$
The first sentence of Theorem 1.6 was stated as a conjecture in an earlier draft of this note, and that conjecture was suggested to the author by Oliver Pechenik. It does not apply in general to weighted graphs. For instance, in the $\overline{p}$-expansion of $K_{21}$ above, the coefficient for $\lambda=3$ is $-1$ while $|\lambda|-\ell(\lambda)=3-1=2$ is even. The statement about the alternating signs of the $a_{W}(k)$’s also does not apply to unweighted graphs. In particular, we will show in §3 that for $K_{21}$ and $W=V(K_{21}),$ $a_{W}(k)$ is positive for $k=2^{j}$, negative for $k=3\cdot 2^{j},$ and 0 otherwise.
The remainder of this note is organized as follows. In §2, we define some necessary terminology. In §3, we revisit the examples above to illustrate Theorem 1.1 and Propositions 1.3 and 1.4. Then in §4, we prove Theorem 1.1, Corollary 1.2, Propositions 1.3 and 1.4, and Corollary 1.5.
@@ 2 Background
A graph $G$ is a set $V$ of vertices together with a set $E$ of edges, where each edge is an unordered pair of vertices. If $uv$ is an edge, we say that the vertices $u$ and $v$ are adjacent. A weighted graph $(G,\omega)$ is a graph $G$ together with a weight function $\omega:V\to\mathbb{Z}_{>0}$ that assigns a positive integer to each vertex. An independent set or stable set is a subset of $V$ in which no two vertices are adjacent. For $W\subseteq V,$ the induced subgraph $G|_{W}$ with vertex set $v$ is the graph with vertex set $V(G|_{W}):=W$ and edge set $E(G|_{W}):=\{uv\in E(G):u,v\in W\}.$ That is, the edges of $G|_{W}$ are precisely the edges of $G$ with both endpoints in $W.$
A symmetric function $f(x_{1},x_{2},\dots)$ is a function that stays the same under any permutation of the variables, i.e. $f(x_{1},x_{2},\dots)=f(x_{\sigma(1)},x_{\sigma(2)},\dots)$ for any permutation $\sigma$ of $\mathbb{Z}_{>0}.$ A proper coloring of $V$ is a function $\alpha:V\to\mathbb{Z}_{>0}$ that assigns a positive integer valued color to each vertex such that $\alpha(u)\neq\alpha(v)$ whenever $uv\in E(G).$ The chromatic symmetric function is
$X_{G}:=\sum_{\alpha}\prod_{v\in V}x_{\alpha(v)}$
@@
A partition $\lambda=\lambda_{1}\lambda_{2}\dots\lambda_{k}$ is a nonincreasing sequence of positive integers $\lambda_{1}\geq\lambda_{2}\geq\dots\geq\lambda_{k},$ and $\lambda_{1},\dots,\lambda_{k}$ are its parts. We write $|\lambda|:=\lambda_{1}+\dots+\lambda_{k}$ for the size of $\lambda$ and $\ell(\lambda)=k$ for its length or number of parts. If $n=|\lambda|,$ we say that $\lambda$ is a partition of $\boldsymbol{n}$ and write $\lambda\vdash n.$ We can alternatively write $\lambda=\ell^{i_{\ell}}(\ell-1)^{i_{\ell-1}}\dots 2^{i_{2}}1^{i_{1}}$ to denote the partition with $i_{j}$ parts of size $j$ for each $j=1,2,\dots,\ell-1,\ell.$ For a partition $\lambda=\lambda_{1}\lambda_{2}\dots\lambda_{k},$ the power sum symmetric function $p_{\lambda}$ is
$p_{\lambda}:=\prod_{i=1}^{k}\left(\sum_{j\geq 1}x_{j}^{\lambda_{i}}\right)=(x_{1}^{\lambda_{1}}+x_{2}^{\lambda_{1}}+\dots)(x_{1}^{\lambda_{2}}+x_{2}^{\lambda_{2}}+\dots)\dots(x_{1}^{\lambda_{k}}+x_{2}^{\lambda_{k}}+\dots).$
A proper set coloring is a function $\alpha:V\to 2^{\mathbb{Z}_{>0}}\backslash\{\varnothing\}$ that assigns a nonempty set of positive integer valued colors to each vertex such that adjacent vertices have nonoverlapping color sets, i.e. $\alpha(u)\cap\alpha(v)=\varnothing$ whenever $uv\in E(G).$ For a weighted graph $(G,\omega)$, the Kromatic symmetric function is
$\overline{X}_{(G,\omega)}:=\sum_{\alpha}\prod_{v\in V}\prod_{i\in\alpha(v)}x_{i}^{\omega(v)}.$
@@
3 Examples
@@
The other subsets of $V$ are $\{v\},$$\{w\},$ and $\varnothing$.∅ . In these cases, we get:
@@
Putting this together, Theorem 1.1 gives
@@
For the seemingly simpler case of a single unweighted edge [Uncaptioned image], there is not such a simple formula for the coefficients:
Example 3.2.
@@
Thus, for $W=V,$ the sequence of exponents $a_{W}(k)$ begins
$a_{W}(1)=2,\ \ a_{W}(2)=-1,\ \ a_{W}(3)=2,\ \ a_{W}(4)=-4,\ \ a_{W}(5)=6,\dots$
These numbers match sequence A038067 in the Online Encyclopedia of Integer Sequences [oeis]. They can be computed recursively using either Proposition 1.3 or Proposition 1.4:
  • To use Proposition 1.3, we first note that
    $\log(I_{G}(t))=\log(1+2t)$ $=2t-\frac{(2t)^{2}}{2}+\frac{(2t)^{3}}{3}-\frac{(2t)^{4}}{4}+\frac{(2t)^{5}}{5}-\cdots$
    $=2t-2t^{2}+\frac{8t^{3}}{3}-4t^{4}+\frac{32t^{5}}{5}-\cdots.$
    Thus we get
    $a_{W}(1)$ $=[t]\log(I_{G}(t))=2,$
    $a_{W}(2)$ $=[t^{2}]\log(I_{G}(t))+\frac{(-1)^{2/1}\cdot 1\cdot a_{W}(1)}{2}=-2+1=-1,$
    $a_{W}(3)$ $=[t^{3}]\log(I_{G}(t))+\frac{(-1)^{3/1}\cdot 1\cdot a_{W}(1)}{3}=\frac{8}{3}-\frac{2}{3}=2,$
    $a_{W}(4)$ $=[t^{4}]\log(I_{G}(t))+\frac{(-1)^{4/2}\cdot 2\cdot a_{W}(2)}{4}+\frac{(-1)^{4/1}\cdot 1\cdot a_{W}(1)}{4}=-4-\frac{1}{2}+\frac{1}{2}=-4,$
    $a_{W}(5)$ $=[t^{5}]\log(I_{G}(t))+\frac{(-1)^{5/1}\cdot 1\cdot a_{W}(1)}{5}=\frac{32}{5}-\frac{2}{5}=6.$
  • Alternatively, to use Proposition 1.4, we list the partitions $k$ for each $k=1,2,3,4,5$:
    • $\boldsymbol{k=1:}$ 1,
    • $\boldsymbol{k=2:}$ 2, $1^{2}$,
    • $\boldsymbol{k=3:}$ 3, 21, $1^{3}$,
    • $\boldsymbol{k=4:}$ 4, 31, $2^{2}$, $21^{2},$ $1^{4}$,
    • $\boldsymbol{k=5:}$ 5, 41, 32, $31^{2}$, $2^{2}1,$ $21^{3},$ $1^{5}.$
    Then we apply Proposition 1.4 to compute
    $a_{W}(1)$ $=[t]I_{G}(t)=2,$
    $a_{W}(2)$ $=[t^{2}]I_{G}(t)-\binom{a_{W}(1)}{2}=0-\binom{2}{2}=0-1=-1,$
    $a_{W}(3)$ $=[t^{3}]I_{G}(t)-\binom{a_{W}(2)}{1}\binom{a_{W}(1)}{1}-\binom{a_{W}(1)}{3}$
    $=0-\binom{-1}{1}\binom{2}{1}-\binom{2}{3}$
    $=0-(-1)(2)-0=2,$
    $a_{W}(4)$ $=[t^{4}]I_{G}(t)-\binom{a_{W}(3)}{1}\binom{a_{W}(1)}{1}-\binom{a_{W}(2)}{2}-\binom{a_{W}(2)}{1}\binom{a_{W}(1)}{2}-\binom{a_{W}(1)}{4}$
    $=0-\binom{2}{1}\binom{2}{1}-\binom{-1}{2}-\binom{-1}{1}\binom{2}{2}-\binom{2}{4}$
    $=0-(2)(2)-1-(-1)(1)-0=-4,$
    $a_{W}(5)$ $=[t^{5}]I_{G}(t)-\binom{a_{W}(4)}{1}\binom{a_{W}(1)}{1}-\binom{a_{W}(3)}{1}\binom{a_{W}(2)}{1}-\binom{a_{W}(3)}{1}\binom{a_{W}(1)}{2}$
    $\hskip 42.67912pt\ \ -\binom{a_{W}(2)}{2}\binom{a_{W}(1)}{1}-\binom{a_{W}(2)}{1}\binom{a_{W}(1)}{3}-\binom{a_{W}(1)}{5}$
    $=0-\binom{-4}{1}\binom{2}{1}-\binom{2}{1}\binom{-1}{1}-\binom{2}{1}\binom{2}{2}-\binom{-1}{2}\binom{2}{1}-\binom{-1}{1}\binom{2}{3}-\binom{2}{5}$
    $=0-(-4)(2)-(2)(-1)-(2)(1)-(1)(2)-(-1)(0)-0=8+2-2-2=6.$
Both methods give us the same exponents $a_{W}(1)=2,\ a_{W}(2)=-1,\ a_{W}(3)=2,\ a_{W}(4)=-4,\ a_{W}(5)=6$ that we listed previously.
To finish computing the $\overline{p}$-expansion of $\overline{X}_{G}$, we note that the other stable sets are $\{v\},$$\{w\}$, and $\varnothing$,∅ , which have independence polynomials
@@
Putting this together, Theorem 1.1 gives
@@
matching the expansion from §1.